### v1

# Dynamic Context Adaptation for Consistent Role-Playing Agents with Retrieval-Augmented Generations

Abstract
The burgeoning prominence of role-playing agents (RPAs) with large language models (LLMs) is driven by their capacity to model complex social interactions and human-like behaviors in controlled, reproducible environments. However, collecting utterances from a specific character or individual is costly and presents substantial challenges for continuously updating character attributes and model parameters, especially for daily streamers or VTubers whose persona information frequently changes. Although retrieval-augmented generation (RAG) can alleviate this problem, there are three underlying limitations. (i) Previous RAG methods truncate each character’s persona paragraph to a fixed length, regardless of hierarchical contexts and varying lengths across characters, which results in responses with lower persona consistency. (ii) If a persona does not contain knowledge relevant to a given query, RAG-based RPAs are prone to hallucination, making it challenging to generate accurate responses. (iii) Existing role-playing datasets are composed of dialogues, and there is no dataset designed for efficiently constructing and evaluating RPAs based on RAG. To address the aforementioned challenges, we propose Amadeus, which is composed of Adaptive Context-aware Text Splitter (ACTS), Guided Selection (GS), and Attribute Extractor (AE). ACTS finds an optimal chunk length and hierarchical contexts for each character. AE identifies a character’s general attributes from the chunks retrieved by GS and uses these attributes as a final context to maintain robust persona consistency even when answering out-of-knowledge questions. To facilitate the development and evaluation of RAG-based RPAs, we construct CharacterRAG, a role-playing dataset that consists of persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs. We find that our framework effectively models not only the knowledge possessed by characters, but also various attributes such as personality.
Introduction
Large language models with long-context capabilities are engineered to manage lengthy input sequences, allowing them to interpret and utilize extended contextual information. (OpenAI 2025; Qwen et al. 2025; Gemini Team 2025, 2024). Although LLMs exhibit enhanced abilities in understanding extended contexts, they still face significant challenges when handling tasks involving genuinely long contexts (Li et al. 2024a). Furthermore, utilizing all relevant information from long-context models to answer each query can be computationally expensive (Li et al. 2024b).
Retrieval-augmented generation (RAG) cost-efficiently mitigates factual inaccuracies and hallucinations in responding to knowledge-intensive queries by integrating external retrieval mechanisms that provide accurate and up-to-date supporting information (Gao et al. 2023; Huang et al. 2025). However, despite these advantages of RAG, there has been little research on RAG-based role-playing agents (RPAs). Moreover, existing role-playing datasets are composed exclusively of dialogues involving characters that are difficult to collect, and there is no dataset designed for building and evaluating RAG-based RPAs.
In this paper, we examine the challenges inherent in RAG-based role-playing and propose approaches to mitigate these challenges. In real-world applications, users and RPAs frequently engage in conversations on topics that extend beyond the knowledge defined in the character’s persona. However, we observe that the existing RAG method tends to excessively utilize chunks that are less relevant to the question when the question is not explicitly answered by the available knowledge (Figure 1). To address this issue and maintain persona consistency, we first introduce an Adaptive Context-aware Text Splitter (ACTS), which segments each character’s persona using optimal chunk length and overlap, thereby preserving context across chunks while simultaneously incorporating hierarchical context information within each chunk. Then, the Guided Selection (GS) retrieves appropriate chunks in order to infer information relevant to the question from the character’s persona such as previous actions and behaviors. Finally, the Attribute Extractor (AE) identifies the general attributes of the character from the retrieved chunks, encouraging the RPA to respond in a manner consistent with that character. Note that if GS fails to find a chunk corresponding to the question, AE extracts the character’s attributes using the chunks with the highest similarity scores, which enables consistent responses.
We conduct extensive experiments to examine the factors influencing the performance of RAG-based role-playing, utilizing various settings with 15 fictional character RPAs, the manually constructed CharacterRAG dataset, and multiple psychological questionnaire datasets. Results demonstrate that our framework reveals new possibilities for RAG-based RPAs. We summarize our contributions as follows:
  • We propose AMADEUS, a RAG-based RPA framework that not only elicits information related to a character, but also maintains persona consistency even when responding to queries beyond its explicit knowledge.
  • We manually construct CharacterRAG, a role-playing dataset for implementing and evaluating RAG-based RPAs comprising persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs.
  • We systematically investigate and uncover key considerations for building RAG-based RPAs through extensive experiments performed in a range of settings.
[…]
Related Work
Role‑Playing Agents. Early persona‑grounded dialogue work such as Persona‑Chat demonstrated that conditioning dialogue models on explicit character descriptions improves conversational coherence and engagement (Zhang et al. 2018). With the advent of LLMs, researchers have pursued finer‑grained persona consistency. Ji et al. (2025) propose a Persona‑aware Contrastive Learning framework that automatically aligns generations with the target character. CharacterGPT (Park, Park, and Lim 2025) proposes a framework that implements four fictional characters through eight attributes. DITTO (Lu et al. 2024) leverages character information to prompt instruction-following LLMs to to engage in role-playing dialogues, treating them as a form of reading comprehension. CharacterGLM (Zhou et al. 2024) introduces an open LLM family that aligns social characters with social traits
Complementary benchmarks soon followed, along with various evaluation methods. RPEval (Boudouri et al. 2025) evaluates emotional understanding, decision‑making, moral alignment, and in‑character fidelity in single‑turn settings, while TimeChara (Ahn et al. 2024) probes point‑in‑time role‑playing and exposes temporal hallucinations in state‑of‑the‑art LLMs. InCharacter (Wang et al. 2024b) adopts an interview-driven methodology to precisely assess RPA personalities by eliciting their underlying mindsets and behaviors. Wang et al. (2025) provides a dynamic, multi-agent virtual environment specifically designed to draw out nuanced, human-like behaviors from large language models during role-playing assessments.
However, current RPAs still truncate persona text, hallucinate beyond retrieved knowledge, and lack a benchmark for incremental updates—gaps our framework fills. In this paper, we propose a novel RAG-based role-playing framework that facilitates the addition and deletion of information.
Retrieval‑Augmented Generation (RAG). RAG couples non-parametric memory with an LLM to mitigate hallucination and stale knowledge (Lewis et al. 2020). Fixed‑length chunking can either omit critical context or waste prompt budget. Recent analyses quantify this trade‑off. Bhat et al. (2025) systematically evaluate fixed‑size chunking across multiple datasets and embedding models, showing that optimal sizes vary widely with domain and answer locality (Bhat et al. 2025). To move beyond one‑size‑fits‑all heuristics, Zhong et al. (2025) introduce Mix‑of‑Granularity, a router that dynamically selects chunk granularity per query and boosts medical QA accuracy (Zhong et al. 2025). Evaluation toolkits are likewise evolving. RAGEval generates scenario‑specific documents, QA pairs, and proposes Completeness, Hallucination, and Irrelevance metrics for rigorous RAG assessment (Zhu et al. 2024). Nevertheless, no existing benchmark explicitly targets role‑playing under RAG, and prior work still assumes personas are short, knowledge‑dense snippets. We mitigate this discrepancy with a novel text splitter that tailors chunk lengths and hierarchical context to each character. We also introduce CharacterRAG, the first dataset designed for building and evaluating RAG‑based role‑playing agents across 15 fictional personas.
Figure 2: An overview of our framework. The Adaptive Context-aware Text Splitter (ACTS) divides a persona into meaningful chunks, selecting the optimal chunk length and overlap for each character. Each chunk is further enhanced with hierarchical context, aiming to minimize information loss about the character. To answer questions beyond the character’s pre-existing knowledge, the Guided Selection (GS) module identifies chunks related to the relevant knowledge, while the Attribute Extractor (AE) extracts general attributes of the character from these chunks. Finally, our framework generates vivid responses by treating the chunks obtained from the RAG as knowledge and the outputs from GS and AE as the character’s attributes.
Task Formulation
Given user query $u$, RAG-based RPAs can be formulated as:
$\mathcal{R}=f(u,\mathcal{D}_{p})$ (1)
, where $\mathcal{D}_{p}$ is a character’s persona, and $f$ is a RPA. A text splitter $g$ divides the persona into $n$ chunks as follows:
$g(\mathcal{D}_{p})=\{c_{1},c_{2},...,c_{n}\}$ (2)
Each chunk $c_{i}$ contains the character’s attributes such as relationships and backstory. The objective of $f$ is to vividly embody the character and generate response $\mathcal{R}$ to $u$ while maintaining persona consistency. However, previous RAG methods (Guu et al. 2020; Guo et al. 2024; Yang et al. 2024; Shukla et al. 2025; Wang, Leung, and Shen 2025) truncate each character’s persona paragraph to a fixed length, regardless of the varying lengths across characters, which results in hallucinations or responses with lower persona consistency. In addition, if the persona does not contain knowledge relevant to the given query, RAG-based role-playing methods are prone to hallucination, making it challenging to generate accurate responses.
Figure 3: Statistics of CharacterRAG Dataset. CharacterRAG consists of persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs.
Method
As shown in Figure 2, our framework consists of (i) Adaptive Context-aware Text Splitter (ACTS), (ii) Guided Selection (GS), and (iii) Attribute Extractor (AE), in order to build a realistic role-playing agent based on RAG.
Adaptive Context-aware Text Splitter
Given the character’s persona $\mathcal{D}_{p}$, ACTS aims to preserve the context between chunks, as well as the information about corresponding subsections within the persona for each chunk, that is, the hierarchical context. To this end, ACTS first finds the maximum length of the paragraphs that constitute the persona as follows:
$l_{\text{max}}=\varphi(p_{1},p_{2},...,p_{l})$ (3)
, where $\varphi$ denotes a length-calculating function. Then, ACTS sets the overlap length of the text splitter to half of $l_{\text{max}}$ (i.e., $l_{\text{o}}=l_{\text{max}}/2$). Note that the reason for setting the overlap length sufficiently large is to minimize information loss, as the context between pieces of information contained in each chunk is indispensable in RAG-based role-playing. Finally, $\mathcal{D}_{p}$ is segmented using $l_{\text{max}}$ and $l_{\text{o}}$ and hierarchical context $\mathcal{H}_{i}$ is appended to each chunk in order to preserve information such as descriptions of the corresponding characters and situational context at each point in the narrative:
$\text{ACTS}(\mathcal{D}_{p},\mathcal{H},l_{\text{max}},l_{\text{o}})=\{\hat{c}_{1},\hat{c}_{2},...,\hat{c}_{m}\}$ (4)
$\hat{c}_{i}=c_{i}+\mathcal{H}_{i}$ (5)
Guided Selection
Algorithm 1 Guided Selection (GS)
Input: User query $u$; Set of knowledge chunks $\mathcal{C}$; Maximum number of search iterations $N$; Slot size $K$
Output: Selected chunk set $S$
1: Initialize slot $S=\emptyset$
2: Sort chunks $\mathcal{C}$ in descending order according to similarity to $u$, obtaining $\mathcal{C}_{\text{sorted}}$
3: Set iteration counter $t\leftarrow 0$
4:for each chunk $c$ in $\mathcal{C}_{\text{sorted}}$ do
5:  if $t\geq N$ or $|S|\geq K$ then
6:   break
7:  end if
8:  With an LLM, determine if chunk $c$ contains information from which the character’s attributes can be inferred regarding $u$
9:  if the LLM returns True then
10:   Add $c$ to slot $S$
11:  end if
12:  $t\leftarrow t+1$
13:end for
14:if $|S|=0$ then
[…]
16:end if
17:return $S$
While RAG has demonstrated significant potential in improving factual correctness of LLMs, existing RAG methods tend to generate uninformative responses (e.g., I’m sorry, but I don’t have specific information.) for questions outside their knowledge base (Guo et al. 2024; Shukla et al. 2025; Wang, Leung, and Shen 2025), or excessively and repetitively use irrelevant chunks that are not pertinent to the given query.
In this paper, GS focuses on selecting appropriate chunks to generate natural and persona-consistent responses. GS is composed of three stages. First, we iterate over the chunks, which are sorted in descending order of similarity to the user query $u$, and employ an LLM to determine whether it is possible to infer the corresponding character’s attributes from each chunk for the $u$. Second, the chunks selected in the previous step are added to the slot, and the iteration terminates when the slot is full. Finally, if the slot remains empty after the maximum number of search iterations, the $K$ chunks with the highest similarity to the query is returned.
GS is effective in identifying chunks containing information that can be inferred from a character’s actions, such as beliefs or personality traits, which are not explicitly stated in the knowledge base and therefore are difficult to retrieve through direct search. For example, if there is no explicit knowledge corresponding to the query ”My living and work spaces are clean and organized” but the information ”Megumin is diligent.” is available, RPA can use this information to infer the characteristics of Megumin and generate an appropriate response. GS is summarized in Algorithm 1.
Attribute Extractor
Inspired by the observations that incorporating character’s attributes can lead to more realistic responses (Park, Park, and Lim 2025; Chen et al. 2025), AE considers two attributes: Belief and Value, and Psychological Traits. Beliefs and values are fundamental principles and ideological orientations that inform and influence a character’s viewpoints and choices. On the other hand, psychological traits refer to characteristics related to personality, emotional states, personal interests, and cognitive tendencies. AE extracts attributes of a character from the chunks generated as a result of GS, and dynamically exploits them as context.
Experiments
Setup
Dataset. We construct the CharacterRAG dataset, which consists of 15 fictional characters, to leverage a RAG-based role-playing framework. CharacterRAG is a high-quality, role-playing dataset in which all external information about works featuring characters that could affect persona consistency has been manually removed, and each persona document has been directly reconstructed from the perspective of each character by human annotators [^1] CharacterRAG dataset is sourced from Namuwiki and is based on Korean data: https://namu.wiki/. For instance, any information speculated from the perspective of editors rather than the characters themselves, as well as information such as character popularity polls that may disrupt role-playing, is excluded. CharacterRAG consists of 15 distinct fictional characters, 976K written char- acters, and 450 question–answer pairs. Six attributes [^2] Unlike the other attributes, Belief and Value and Psychological Traits directly impact the behavior of a character; therefore, AE utilizes information from these two attributes. define each character’s persona and the corresponding QA pairs (Chen et al. 2025):
  • Activity: A documented history comprising prior activities, behaviors, and interactions, encompassing elements such as backstory and schedules.
  • Belief and Value: The foundational tenets, dispositions, and ideological orientations that inform and guide a character’s viewpoints and decision-making processes (e.g., beliefs and attitudes).
  • Demographic Information: Information that can identify a character, including but not limited to their name, age, educational background, professional history, and geographic location.
  • Psychological Traits: Attributes associated with personality traits, emotional states, preferences, and patterns of cognitive behavior.
  • Skill and Expertise: The extent of understanding, skillfulness, and competence regarding particular fields or technologies.
  • Social Relationships: The characteristics and processes of social interactions, encompassing individuals’ roles, relational ties, and patterns of communication.
[…]
$\sum\mu$ $\sum\sigma^{2}$ $\sum\mu$ $\sum\sigma^{2}$ $\sum\mu$ $\sum\sigma^{2}$
Vanilla 6.4325 0.1026 8.3306 0.1557 12.3136 0.0071
ATS 6.7007 0.0884 8.4718 0.1281 12.2336 0.0070
ACTS (Ours) 6.8575 0.0784 8.6226 0.1179 12.3240 0.0063
Table 1: Distribution of Similarity Scores. Vanilla refers to the method of splitting chunks with a fixed length and overlap, ATS refers to dividing a persona for each character with optimal segment length and overlap, without hierarchical context.
$\mu$ $\sigma$ Mdn Cronbach’s alpha
BFI 3.890 0.965 4.100 0.845
MBTI 3.870 0.911 3.933 0.830
Table 2: Human Evaluation. We conduct human evaluation using a 5-point Likert scale to verify whether inferring character attributes with the Attribute Extractor (AE) from chunks extracted by Guided Selection (GS) is reasonable. Note that $S$ represents the results of all human evaluators, $\mu$ is $\mathbb{E}(\mathbb{E}(S))$, $\sigma$ is $\mathbb{E}(\sigma(S))$, and Mdn is $\mathbb{E}(Mdn(S)).$
Chunk Usage Rate GS Failure
CharacterRAG 67.56% 25.11%
MBTI 43.84% 15.33%
Table 3: Average Chunk Usage Rate and GS Failure. We present 30 CharacterRAG questions and 60 MBTI questions to 15 characters, and measure the average chunk usage rate and GS failure.
Each section of the character’s persona contains subsections, preserving the hierarchical context of each chunk (e.g., ”Frieren’s activities in the story, the journey to find Aureole, the First-Class Mage Exam”). Furthermore, Each QA pair consists of a question and corresponding answer derived from the character’s knowledge, pertaining to one of the six attributes manually constructed for each character. Detailed statistics of CharacterRAG are shown in Figure 3.
[…]
Settings. We implement Guided Selection (GS) and Attribute Extractor (AE) using GPT-4.1 (”gpt-4.1-2025-04-14”). We apply the ACTS to the conventional Naive RAG. The maximum number of search iterations $N$ is 30, and the slot size $K$ is set to 2. Note that the speed at which GS identifies and retrieves relevant chunks for answering is much faster and more cost-efficient than that of previous graph-based or web search-based RAG methods (Shukla et al. 2025; Guo et al. 2024; Yan et al. 2024).
Evaluation Protocols
Tasks. We use 450 QA pairs from the CharacterRAG dataset to verify whether the RPA sufficiently leverages each character’s knowledge. We also use 60 MBTI [^3] https://www.16personalities.com/ questions and 120 BFI (Barrick and Mount 1991) questions to investigate whether each character can appropriately respond to questions for which they do not have explicit prior knowledge. Since it is not possible to construct QA pairs for questions outside the scope of a character’s knowledge, we instead conduct an interview-based assessment (Wang et al. 2024b) for each character and compare the results to psychological test outcomes for the character, as determined by thousands of actual participants’ votes [^4] https://www.personality-database.com/.
Metrics. We design three metrics to extensively evaluate the role-playing capabilities of RAG-based RPAs: (i) ACC measures whether the character’s response contains the correct answer or not. (ii) ACCL is a score assigned by the LLM, ranging from 1 to 10, that evaluates how well the character’s response reflects the correct answer. (iii) Hallucination Score (HS) evaluates the degree of hallucination in the model’s response given a query, the relevant chunks, and the ground-truth answer, on a scale from 1 to 10. Specifically, HS is assigned close to 1 when the response faithfully reflects only the facts contained in the chunks or answer without distortion or addition, indicating minimal hallucination.
CRAG LightRAG AMADEUS (Ours) GT
Anya Forger INTJ (-3) INFJ (-2) ENFP (0) ENFP
Chika Fujiwara ENFP (0) INTP (-2) ENFP (0) ENFP
Edward Elric ISTJ (-3) INTP (-1) INFP (-2) ENTP
Frieren INFP (-1) INTP (0) INTP (0) INTP
Hitori Gotoh ISTJ (-2) ENFP (-1) INFP (0) INFP
Light Yagami INTJ (-1) INTJ (-1) INTJ (-1) ENTJ
Māo Māo INTJ (-1) INTJ (-1) ISTP (-1) INTP
Megumin INFP (0) INFP (0) ISFP (-1) INFP
Mikoto Misaka ISFP (-4) ENTJ (0) INFJ (-2) ENTJ
Nina Iseri ISFP (0) ENFJ (-3) INFP (-1) ISFP
Saitama ISTP (0) INTP (-1) ISTP (0) ISTP
Son Goku ENFJ (-2) INTP (-3) ESFP (0) ESFP
Tanjiro Kamado ENFJ (0) INTP (-3) ENFJ (0) ENFJ
Tobio Kageyama ENTJ(-2) INFJ (-1) ISTJ (0) ISTJ
Yui Hirasawa ENFP (0) INTP (-2) ESFP (-1) ENFP
Accuracy 68.33% 65.00% 85.00% -
Avg F1-Score 0.6448 0.5344 0.8244 -
Table 4: Predicted MBTI Types per Character. The number in parentheses indicates the number of times the ground-truth (GT) type of each character was not correctly identified (using the GPT-4.1 setting).
Experimental Results
[…]
CRAG LightRAG AMADEUS (Ours) GT
Anya Forger SLOAI (-2) RCUEN (-3) SLUEI (-2) SCUAI
Chika Fujiwara SCUAI (0) RLUEN (-4) SCOAI (-1) SCUAI
Edward Elric SLOEI (-1) RCUAN (-4) SLOEI (-1) SLUEI
Frieren RCUAI (-1) SLUEN (-3) RCUAI (-1) RCUEI
Hitori Gotoh RLUAI (0) RCUEN (-3) RLUAI (0) RLUAI
Light Yagami SCOEI (-1) RCUAN (-3) SCOEI (-1) RCOEI
Māo Māo RCOAN (-2) SLUAN (-5) RCOEI (0) RCOEI
Megumin SCUAI (-2) RLUEN (-2) SLOEI (-1) SLUEI
Mikoto Misaka SLOAI (-3) RCUEN (-2) SLOAI (-3) RCOEI
Nina Iseri RLUAI (-1) RCUEN (-2) SLUEI (-1) RLUEI
Saitama RCOAN (-1) SCOAI (-3) RCUAN (0) RCUAN
Son Goku SCUAI (-1) RLUEI (-4) SCOAI (-2) SCUAN
Tanjiro Kamado SCOAI (0) RCUAN (-3) SCOAI (0) SCOAI
Tobio Kageyama SLOAN (-2) RCUAI (-4) RCOEN (-1) RLOEN
Yui Hirasawa SLUAI (-1) RCOEN (-4) SCUAI (0) SCUAI
Accuracy 76.00% 34.67% 81.33% -
Avg F1-Score 0.7313 0.2774 0.7986 -
Table 5: Predicted Big 5 SLOAN Types per Character. The number in parentheses indicates the number of times the ground-truth (GT) type of each character was not correctly identified (using the GPT-4.1 setting).
Extracting a character’s attributes from selected text chunks is reliable. We investigate the reasonableness of inferring character attributes with the Attribute Extractor (AE) from chunks extracted via Guided Selection (GS) by conducting human evaluation using a 5-point Likert scale. To this end, we invite 11 human evaluators and each evaluator is asked to score 60 randomly selected samples. Each sample consists of pairs of chunks selected from GS and attributes extracted through AE, for 30 BFI questions and 30 MBTI questions that are not included in the character’s knowledge.
In Table 2, we find that the means $\mu$ is close to 4, with small standard deviations $\sigma$. It demonstrates that the outputs of GS and AE are reliable and trustworthy, even from a human evaluative perspective. We also measure Cronbach’s alpha (Cronbach 1951) to evaluate internal consistency among the human evaluators. We find that the Cronbach’s alpha values are 0.845 and 0.830, both exceeding the commonly accepted threshold of 0.7 for acceptable reliability. Since values above 0.8 are generally interpreted as indicating a high level of internal consistency, experimental results in Table 2 can be considered highly trustworthy.
Selecting appropriate chunks increases the utilization of a character’s persona. Table 3 shows the usage rate of persona chunks and the failure rate of GS when each each of the 15 characters is presented with questions that are both within and outside their knowledge. Note that, despite the limited number of queries in CharacterRAG, the Chunk Usage Rate still reaches 67.56%. The findings demonstrate that GS serves as an effective tool for enabling the use of character personas. At the same time, it can be observed that the 450 questions in CharacterRAG are designed to capture the core information of each character.
Meanwhile, as shown in Table 3 and Figure 1, GS substantially increases the chunk utilization rate for the 60 MBTI questions, from 34.93% to 43.84%. The proportion of cases where the GS fails to find an appropriate chunk, i.e., GS Failure, is 25.11% and 15.33%, respectively, which demonstrates that although GS effectively utilizes the character’s persona, there is still room for further improvement. Nevertheless, despite this shortcoming, our framework offers markedly enhanced speed, cost-effectiveness, and stable performance compared to earlier RAG approaches utilizing graph structures or web search mechanisms. (Shukla et al. 2025; Guo et al. 2024; Yan et al. 2024).
Method Task GPT-4.1 Gemma3-27B Qwen3-32B
[…]
w/o RAG CharacterRAG 49.56% 6.79 - 27.56% 5.33 - 18.89% 4.35 -
Naive RAG CharacterRAG 91.33% 9.23 3.13 86.44% 8.85 3.27 78.44% 8.49 5.05
LightRAG CharacterRAG 48.00% 6.06 - 69.56% 8.17 - 68.67% 8.20 -
CRAG CharacterRAG 70.00% 8.26 3.21 57.78% 7.57 4.09 28.67% 5.24 8.68
AMADEUS (Ours) CharacterRAG 92.67% 9.26 2.89 88.00% 8.92 3.26 78.89% 8.63 4.66
Naive RAG MBTI - - 2.69 - - 2.53 - - 2.33
CRAG MBTI - - 2.38 - - 2.91 - - 1.80
AMADEUS (Ours) MBTI - - 2.05 - - 2.02 - - 2.04
Naive RAG BFI - - 2.74 - - 2.52 - - 2.42
CRAG BFI - - 2.26 - - 2.75 - - 1.96
AMADEUS (Ours) BFI - - 1.94 - - 1.99 - - 2.03
Table 6: Evaluation of role-playing capabilities on CharacterRAG, MBTI, and BFI tasks. Higher values of ACC (%) and ACCL (1-10) correspond to better performance, whereas lower values of HS (1-10) are preferable.
Graph-based RAG and web search-based RAG are unsuitable for role-playing. One of the major challenges in retrieval-based role-playing is that, when a RPA receives questions involving knowledge outside a character’s persona, it tends to either overuse irrelevant chunks (Figure 1) or generate uninformative responses (Guo et al. 2024; Shukla et al. 2025; Wang, Leung, and Shen 2025).
To investigate whether RAG-based RPAs can handle this problem, we conduct extensive experiments in which we ask 15 characters 60 MBTI questions and 120 BFI questions each, and and evaluate their ability to accurately infer the characters’ personality types. Table 4 and Table 5 show predicted MBTI types and Big 5 SLOAN types per character. Our framework maintains persona consistency even when answering questions that are not explicitly specified in each character’s persona in both MBTI and BFI settings. Note that the performance gap of CRAG is significant between the two settings. We assume that questions requiring analogical reasoning are difficult to solve even with web search and that the search results may contain non-negligible noise. On the other hand, LightRAG exhibits the lowest performance, which shows that graph-based RAG methods are not well suited for RPA applications due to the high cost of graph construction, the difficulty in adding or removing new knowledge, and challenges in maintaining persona consistency. While we did not perform a direct comparative experiment, we observed that GraphRAG (Shukla et al. 2025) suffers from similar problems. Specifically, applying graph-based RAG to RPA is challenging in cases where the persona frequently changes, such as with daily streamers or VTubers.
CharacterRAG dataset serves as a valuable resource for the construction and evaluation of RAG-based RPAs. To investigate the factors influencing the performance of role-playing, we conduct a comprehensive interview-based assessments on the generalization capabilities of models with various LLMs and RAG techniques. Table 6 presents how the ability to accurately answer questions related to the character’s knowledge, which is a core aspect of role-playing, varies across the applied methodologies.
We first examine to what extent each LLM possesses background knowledge about the 15 characters in a setting without RAG, and results show that none of the three LLMs are capable of effective role-playing without access to external knowledge. Moreover, we observe that LightRAG, a graph-based RAG, is ill-suited for the storage and retrieval of character knowledge, as it often suffers from issues such as entity ambiguity and uninformative responses. In a similar vein, CRAG exhibits challenges in maintaining role-playing fidelity, which can be attributed to the tendency of web search-based RAG methods to utilize retrieved content that may undermine the consistency of a character’s persona. Indeed, despite leveraging web information, CRAG is able to correctly answer only 6 out of the 30 CharacterRAG questions pertaining to Nina Iseri. In addition, to analyze how a thinking mode of LLMs influences their role-playing capabilities, we employ Qwen 3-32B. Results demonstrate that the thinking mode fails to yield any substantial positive effect on enhancing role-playing performance.
Note that our framework achieves the best performance across all three LLMs. We also find that the Hallucination Score (HS) is the lowest in CharacterRAG setting. These results highlight the importance of preserving the context of split-character personas and effectively leveraging appropriate character attributes in RAG-based RPAs. Furthermore, such elements are especially pronounced in dialogue situations that transcend the scope of the character’s knowledge (Table 4 and Table 5). We believe that our findings demonstrate new possibilities for RAG-based RPAs.
Conclusion
In this work, we addresse critical limitations in building retrieval-augmented, RPAs with LLMs. By introducing a novel framework consisting of an Adaptive Context-aware Text Splitter (ACTS), Guided Selection (GS), and Attribute Extractor (AE), our approach enables robust and consistent simulation of character personas, even when confronted with queries that extend beyond explicit persona knowledge. Through the development of the CharacterRAG dataset, we provide a valuable resource for reproducible evaluation and benchmarking of RAG-based RPAs. Our experimental results demonstrate that the proposed method not only enhances the character’s knowledge representation, but also faithfully models nuanced traits such as personality. We are enthusiastic about the future prospects of RAG-driven role-playing agents, along with the creation of stronger character personas and improved RAG architectures.

### v3

@@
 Abstract
-The burgeoning prominence of role-playing agents (RPAs) with large language models (LLMs) is driven by their capacity to model complex social interactions and human-like behaviors in controlled, reproducible environments.
-However, collecting utterances from a specific character or individual is costly and presents substantial challenges for continuously updating character attributes and model parameters, especially for daily streamers or VTubers whose persona information frequently changes.
-Although retrieval-augmented generation (RAG) can alleviate this problem, there are three underlying limitations.
-(i) Previous RAG methods truncate each character’s persona paragraph to a fixed length, regardless of hierarchical contexts and varying lengths across characters, which results in responses with lower persona consistency.
-(ii) If a persona does not contain knowledge relevant to a given query, RAG-based RPAs are prone to hallucination, making it challenging to generate accurate responses.
-(iii) Existing role-playing datasets are composed of dialogues, and there is no dataset designed for efficiently constructing and evaluating RPAs based on RAG.
-To address the aforementioned challenges, we propose Amadeus, which is composed of Adaptive Context-aware Text Splitter (ACTS), Guided Selection (GS), and Attribute Extractor (AE).
-ACTS finds an optimal chunk length and hierarchical contexts for each character.
-AE identifies a character’s general attributes from the chunks retrieved by GS and uses these attributes as a final context to maintain robust persona consistency even when answering out-of-knowledge questions.
-To facilitate the development and evaluation of RAG-based RPAs, we construct CharacterRAG, a role-playing dataset that consists of persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs.
-We find that our framework effectively models not only the knowledge possessed by characters, but also various attributes such as personality.
-Introduction
-Large language models with long-context capabilities are engineered to manage lengthy input sequences, allowing them to interpret and utilize extended contextual information.
-(OpenAI 2025; Qwen et al.
-2025; Gemini Team 2025, 2024).
-Although LLMs exhibit enhanced abilities in understanding extended contexts, they still face significant challenges when handling tasks involving genuinely long contexts (Li et al.
-2024a).
-Furthermore, utilizing all relevant information from long-context models to answer each query can be computationally expensive (Li et al.
-2024b).
-Retrieval-augmented generation (RAG) cost-efficiently mitigates factual inaccuracies and hallucinations in responding to knowledge-intensive queries by integrating external retrieval mechanisms that provide accurate and up-to-date supporting information (Gao et al.
-2023; Huang et al.
-2025).
+Building role-playing agents (RPAs) that faithfully emulate specific characters remains challenging because collecting character-specific utterances and continually updating model parameters are resource-intensive, making retrieval-augmented generation (RAG) a practical necessity.
+However, despite the importance of RAG, there has been little research on RAG-based RPAs.
+For example, we empirically find that when a persona lacks knowledge relevant to a given query, RAG-based RPAs are prone to hallucination, making it challenging to generate accurate responses.
+In this paper, we propose Amadeus, a training-free framework that can significantly enhance persona consistency even when responding to questions that lie beyond a character’s knowledge.
+In addition, to underpin the development and rigorous evaluation of RAG-based RPAs, we manually construct CharacterRAG, a role-playing dataset that consists of persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs.
+We find that our proposed method effectively models not only the knowledge possessed by characters, but also various attributes such as personality.
+The code and dataset will be made publicly available at our GitHub.
+Dynamic Context Adaptation for Consistent Role-Playing Agents with Retrieval-Augmented Generations
+Jeiyoon Park1, Yongshin Han1, Minseop Kim1, Kisu Yang2, 1SOOP, 2Korea University {naruto, winterfeb, usio}@sooplive.com, willow4@korea.ac.kr https://huggingface.co/datasets/naruto-soop/CharacterRAG
+Figure 1: Comparison of Responses between Naive RAG and Our Method.
+We observe that the existing RAG method tends to excessively utilize chunks that are less relevant to the question when the question is not explicitly answered by the available knowledge.
+1 Introduction
+Large language models with long-context capabilities are engineered to manage lengthy input sequences, allowing them to interpret and utilize extended contextual information (OpenAI, 2025; Qwen et al., 2025; Gemini Team, 2025, 2024).
+Although LLMs exhibit enhanced abilities in understanding extended contexts, they still face significant challenges when handling tasks involving genuinely long contexts (Li et al., 2024a).
+Furthermore, utilizing all relevant information from long-context models to answer each query can be computationally expensive (Li et al., 2024b).
+Retrieval-augmented generation (RAG) cost-efficiently mitigates factual inaccuracies and hallucinations in responding to knowledge-intensive queries by integrating external retrieval mechanisms that provide accurate and up-to-date supporting information (Gao et al., 2023; Huang et al., 2025).
 However, despite these advantages of RAG, there has been little research on RAG-based role-playing agents (RPAs).
 Moreover, existing role-playing datasets are composed exclusively of dialogues involving characters that are difficult to collect, and there is no dataset designed for building and evaluating RAG-based RPAs.
-In this paper, we examine the challenges inherent in RAG-based role-playing and propose approaches to mitigate these challenges.
+We examine the challenges inherent in RAG-based role-playing and propose approaches.
 In real-world applications, users and RPAs frequently engage in conversations on topics that extend beyond the knowledge defined in the character’s persona.
-However, we observe that the existing RAG method tends to excessively utilize chunks that are less relevant to the question when the question is not explicitly answered by the available knowledge (Figure 1).
-To address this issue and maintain persona consistency, we first introduce an Adaptive Context-aware Text Splitter (ACTS), which segments each character’s persona using optimal chunk length and overlap, thereby preserving context across chunks while simultaneously incorporating hierarchical context information within each chunk.
-Then, the Guided Selection (GS) retrieves appropriate chunks in order to infer information relevant to the question from the character’s persona such as previous actions and behaviors.
-Finally, the Attribute Extractor (AE) identifies the general attributes of the character from the retrieved chunks, encouraging the RPA to respond in a manner consistent with that character.
-Note that if GS fails to find a chunk corresponding to the question, AE extracts the character’s attributes using the chunks with the highest similarity scores, which enables consistent responses.
-We conduct extensive experiments to examine the factors influencing the performance of RAG-based role-playing, utilizing various settings with 15 fictional character RPAs, the manually constructed CharacterRAG dataset, and multiple psychological questionnaire datasets.
-Results demonstrate that our framework reveals new possibilities for RAG-based RPAs.
-We summarize our contributions as follows:
-  • We propose AMADEUS, a RAG-based RPA framework that not only elicits information related to a character, but also maintains persona consistency even when responding to queries beyond its explicit knowledge.
-  • We manually construct CharacterRAG, a role-playing dataset for implementing and evaluating RAG-based RPAs comprising persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs.
-  • We systematically investigate and uncover key considerations for building RAG-based RPAs through extensive experiments performed in a range of settings.
+However, we observe that the existing RAG method tends to excessively utilize chunks that are less relevant to the question when the question is not explicitly answered by the available knowledge (Figure 1 and Figure 6).
+To address these challenges, we introduce Amadeus, a training-free framework that can markedly improves persona consistency, even when addressing questions beyond a character’s knowledge.
+Amadeus consists of three substages: Adaptive Context-aware Text Splitter (ACTS), which segments a persona for role-playing, Guided Selection (GS), which retrieves appropriate chunks to infer information relevant to the question from the persona, such as prior actions and behaviors, and Attribute Extractor (AE), which identifies general attributes of the character from the retrieved chunks, thereby encouraging the RPA to respond in a manner consistent with that character.
+To underpin the development and rigorous evaluation of RAG-based RPAs, we manually construct CharacterRAG, a role-playing dataset that consists of persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs.
+We conduct extensive experiments to examine the factors that influence the performance of RAG-based RPAs through interview-based assessments (Wang et al., 2024b; Park et al., 2025; Jiang et al., 2023).
+To investigate whether RPAs can respond appropriately to questions that fall outside their endowed knowledge, we employ multiple psychological questionnaires such as MBTI [^1] https://www.16personalities.com/ and BFI (Barrick and Mount, 1991), following prior studies Sang et al.
+(2022); Wang et al.
+(2024b).
+Results show that our framework opens up new possibilities for RAG-based role-playing.
+2 CharacterRAG
+2.1 Dataset Construction
+We construct the CharacterRAG dataset, which consists of 15 fictional characters, to leverage and evaluate a RAG-based role-playing framework.
+CharacterRAG is a high-quality, role-playing dataset in which all external information about works featuring characters that could affect persona consistency has been manually removed, and each persona document has been directly reconstructed from the perspective of each character by human annotators [^2] CharacterRAG dataset is sourced from Namuwiki and is based on Korean data: https://namu.wiki/.
+For instance, any information speculated from the perspective of editors rather than the characters themselves, as well as information such as character popularity polls that may disrupt role-playing, is excluded.
+CharacterRAG consists of 15 distinct fictional characters, 976K written characters, and 450 question–answer pairs.
+Figure 2: An overview of CharacterRAG Dataset. CharacterRAG consists of persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs.
+2.2 Attributes
+Six commonly used attributes in role‑playing define each character’s persona and the corresponding question–answer pairs (Chen et al., 2025): (1) Activity.
+A documented history comprising prior activities, behaviors, and interactions, encompassing elements such as backstory and schedules.
+(2) Belief and Value.
+The foundational tenets, dispositions, and ideological orientations that inform and guide a character’s viewpoints and decision-making processes (e.g., beliefs and attitudes).
+(3) Demographic Information.
+Information that can identify a character, including but not limited to their name, age, educational background, professional history, and geographic location.
+(4) Psychological Traits.
+Attributes associated with personality traits, emotional states, preferences, and patterns of cognitive behavior.
+(5) Skill and Expertise.
+The extent of understanding, skillfulness, and competence regarding particular fields or technologies.
+(6) Social Relationships.
+The characteristics and processes of social interactions, encompassing individuals’ roles, relational ties, and patterns of communication.
+Each section of the character’s persona contains subsections, preserving the hierarchical information (e.g., "Tanjiro Kamado’s actions in the story").
+Furthermore, Each QA pair consists of a question and corresponding answer derived from the character’s knowledge, pertaining to one of the six attributes manually constructed for each character.
+Statistics and samples of CharacterRAG are shown in Figure 2 and Table 7.
+3 Task Formulation: RAG-based Role-Playing Agents
+Given user query $u$, RAG-based RPAs can be formulated as: $\mathcal{R}=f(u,\mathcal{D}_{p})$, where $\mathcal{D}_{p}$ is a character’s persona, and $f$ is a RPA.
+A text splitter $g$ divides the persona into $n$ chunks as follows:
+$g(\mathcal{D}_{p})=\{c_{1},c_{2},...,c_{n}\}$ (1)
+Each chunk $c_{i}$ contains a character’s knowledge corresponding to attributes in Section 2.2.
+Rather than using the full chunks $\mathcal{C}=\{c_{i}\mid i=1,\dots,n\}$, $f$ takes as input the top $K$ chunks with the highest semantic scores relative to the $u$:
+$\mathcal{C}^{*}=\text{TopK}(\{\text{sim}(u,c_{i})\}_{i=1}^{n}),\quad|\mathcal{C}^{*}|=K$ (2)
+The objective of $f$ is to vividly embody a character and generate response $\mathcal{R}$ to $u$ while maintaining persona consistency: $\mathcal{R}^{*}=f(u,\mathcal{C}^{*})$.
+However, previous RAG methods (Guu et al., 2020; Guo et al., 2024; Yang et al., 2024; Shukla et al., 2025; Wang et al., 2025b) truncate each character’s persona paragraph to a fixed length, regardless of the varying lengths across characters, which results in hallucinations or responses with lower persona consistency.
+Although existing works (LangChain, 2023; Antematter, 2024; Zhong et al., 2025a; Liu et al., 2025) explore optimal chunking strategies, they struggle to capture the contextual similarities across chunks that are crucial for role-playing.
+4 Method
+Amadeus consists of three substages: Adaptive Context-aware Text Splitter (ACTS), Guided Selection (GS), and Attribute Extractor (AE), in order to build realistic RAG-based RPAs (Figure 5).
+4.1 Adaptive Context-aware Text Splitter
+Unlike previous naive semantic chunking methods, Given $\mathcal{D}_{p}$, ACTS aims to preserve intra-level context across chunks and, for each chunk, information about the corresponding subsections of the persona—that is, hierarchical context $\mathcal{H}$.
+For instance, in Figure 2, chunks within ### 4.1.1 must preserve $\mathcal{H}$: "Tanjiro Kamado’s actions in the story (# 4)", "The death of his family (## 4.1)", and "Tanjiro vs.
+the Water Hashira, Giyu Tomiok (### 4.1.1)".
+To this end, ACTS first finds the maximum length of the paragraphs that constitute the persona: $l_{\text{max}}=\varphi(p_{1},p_{2},...,p_{l})$, where $\varphi$ denotes a length-calculating function.
+Then, ACTS sets the overlap length of the text splitter to half of $l_{\text{max}}$ (i.e., $l_{\text{o}}=l_{\text{max}}/2$).
+The reason for setting both chunk length and overlap length sufficiently large is to minimize information loss, as the context between pieces of information contained in each chunk is indispensable in RAG-based role-playing (Figure 8).
+Finally, ACTS recursively retrieves the context at each chunk’s current position in the hierarchy, then segments $\mathcal{D}_{p}$ using $l_{\text{max}}$ and $l_{\text{o}}$, and concatenates the resulting context $\mathcal{H}_{i}$ to each chunk to preserve information such as character descriptions and situational context at each point in the narrative: $\text{ACTS}(\mathcal{D}_{p},\mathcal{H},l_{\text{max}},l_{\text{o}})=\{\hat{c}_{1},\hat{c}_{2},...,\hat{c}_{m}\}$, where $\hat{c}_{i}=[c_{i};\mathcal{H}_{i}]$.
+From a computational standpoint, the extraction of hierarchical context incurs an O(N) runtime cost, making it efficient.
+4.2 Guided Selection
+GS focuses on selecting appropriate chunks to generate natural and persona-consistent responses.
+GS is composed of three stages: (1) We iterate over the chunks, which are sorted in descending order of semantic similarity to the user query $u$, and employ an LLM to determine whether it is possible to infer the corresponding character’s attributes from each chunk for the $u$.
+(2) The chunks selected in the previous step are appended to the slot, and the iteration terminates when the slot is full.
+(3) If the slot remains empty after the maximum number of search iterations, the $K$ chunks with the highest semantic similarity to the query is returned.
+Note that GS is effective in identifying chunks containing information that can be inferred from a character’s actions, such as beliefs or personality traits, which are not explicitly stated in the knowledge base and therefore are difficult to retrieve through direct search.
+For example, even if there is no explicit knowledge corresponding to the query "My living and work spaces are clean and organized", if Megumin’s conscientiousness can be inferred from her behavior depicted in the narrative, RPA can leverage this information to infer the characteristics of Megumin and generate an appropriate response.
+GS is summarized in Algorithm 1 and Figure 3.
+4.3 Attribute Extractor
+Inspired by the observations that incorporating character’s attributes can lead to more realistic responses (Park et al., 2025; Chen et al., 2025), AE considers two attributes: Belief and Value, and Psychological Traits.
+Beliefs and values are fundamental principles and ideological orientations that inform and influence a character’s viewpoints and choices.
+On the other hand, psychological traits refer to characteristics related to personality, emotional states, personal interests, and cognitive tendencies.
+AE extracts attributes of a character from the chunks generated as a result of GS, and exploits them as context.
+Finally, we dynamically construct the context using knowledge retrieved via RAG and attributes extracted by GS and AE, enabling the model to generate vivid responses (Figure 7).
+5 Experiments
+5.1 Setup
 […]
-Related Work
-Role‑Playing Agents.
-Early persona‑grounded dialogue work such as Persona‑Chat demonstrated that conditioning dialogue models on explicit character descriptions improves conversational coherence and engagement (Zhang et al.
-2018).
-With the advent of LLMs, researchers have pursued finer‑grained persona consistency.
-Ji et al.
-(2025) propose a Persona‑aware Contrastive Learning framework that automatically aligns generations with the target character.
-CharacterGPT (Park, Park, and Lim 2025) proposes a framework that implements four fictional characters through eight attributes.
-DITTO (Lu et al.
-2024) leverages character information to prompt instruction-following LLMs to to engage in role-playing dialogues, treating them as a form of reading comprehension.
-CharacterGLM (Zhou et al.
-2024) introduces an open LLM family that aligns social characters with social traits
-Complementary benchmarks soon followed, along with various evaluation methods.
-RPEval (Boudouri et al.
-2025) evaluates emotional understanding, decision‑making, moral alignment, and in‑character fidelity in single‑turn settings, while TimeChara (Ahn et al.
-2024) probes point‑in‑time role‑playing and exposes temporal hallucinations in state‑of‑the‑art LLMs.
-InCharacter (Wang et al.
-2024b) adopts an interview-driven methodology to precisely assess RPA personalities by eliciting their underlying mindsets and behaviors.
-Wang et al.
-(2025) provides a dynamic, multi-agent virtual environment specifically designed to draw out nuanced, human-like behaviors from large language models during role-playing assessments.
-However, current RPAs still truncate persona text, hallucinate beyond retrieved knowledge, and lack a benchmark for incremental updates—gaps our framework fills.
-In this paper, we propose a novel RAG-based role-playing framework that facilitates the addition and deletion of information.
-Retrieval‑Augmented Generation (RAG).
-RAG couples non-parametric memory with an LLM to mitigate hallucination and stale knowledge (Lewis et al.
-2020).
-Fixed‑length chunking can either omit critical context or waste prompt budget.
-Recent analyses quantify this trade‑off.
-Bhat et al.
-(2025) systematically evaluate fixed‑size chunking across multiple datasets and embedding models, showing that optimal sizes vary widely with domain and answer locality (Bhat et al.
-2025).
-To move beyond one‑size‑fits‑all heuristics, Zhong et al.
-(2025) introduce Mix‑of‑Granularity, a router that dynamically selects chunk granularity per query and boosts medical QA accuracy (Zhong et al.
-2025).
-Evaluation toolkits are likewise evolving.
-RAGEval generates scenario‑specific documents, QA pairs, and proposes Completeness, Hallucination, and Irrelevance metrics for rigorous RAG assessment (Zhu et al.
-2024).
-Nevertheless, no existing benchmark explicitly targets role‑playing under RAG, and prior work still assumes personas are short, knowledge‑dense snippets.
-We mitigate this discrepancy with a novel text splitter that tailors chunk lengths and hierarchical context to each character.
-We also introduce CharacterRAG, the first dataset designed for building and evaluating RAG‑based role‑playing agents across 15 fictional personas.
-Figure 2: An overview of our framework.
-The Adaptive Context-aware Text Splitter (ACTS) divides a persona into meaningful chunks, selecting the optimal chunk length and overlap for each character.
-Each chunk is further enhanced with hierarchical context, aiming to minimize information loss about the character.
-To answer questions beyond the character’s pre-existing knowledge, the Guided Selection (GS) module identifies chunks related to the relevant knowledge, while the Attribute Extractor (AE) extracts general attributes of the character from these chunks.
-Finally, our framework generates vivid responses by treating the chunks obtained from the RAG as knowledge and the outputs from GS and AE as the character’s attributes.
-Task Formulation
-Given user query $u$, RAG-based RPAs can be formulated as:
-$\mathcal{R}=f(u,\mathcal{D}_{p})$ (1)
-, where $\mathcal{D}_{p}$ is a character’s persona, and $f$ is a RPA. A text splitter $g$ divides the persona into $n$ chunks as follows:
-$g(\mathcal{D}_{p})=\{c_{1},c_{2},...,c_{n}\}$ (2)
-Each chunk $c_{i}$ contains the character’s attributes such as relationships and backstory.
-The objective of $f$ is to vividly embody the character and generate response $\mathcal{R}$ to $u$ while maintaining persona consistency.
-However, previous RAG methods (Guu et al.
-2020; Guo et al.
-2024; Yang et al.
-2024; Shukla et al.
-2025; Wang, Leung, and Shen 2025) truncate each character’s persona paragraph to a fixed length, regardless of the varying lengths across characters, which results in hallucinations or responses with lower persona consistency.
-In addition, if the persona does not contain knowledge relevant to the given query, RAG-based role-playing methods are prone to hallucination, making it challenging to generate accurate responses.
-Figure 3: Statistics of CharacterRAG Dataset. CharacterRAG consists of persona documents for 15 distinct fictional characters totaling 976K written characters, and 450 question–answer pairs.
-Method
-As shown in Figure 2, our framework consists of (i) Adaptive Context-aware Text Splitter (ACTS), (ii) Guided Selection (GS), and (iii) Attribute Extractor (AE), in order to build a realistic role-playing agent based on RAG.
-Adaptive Context-aware Text Splitter
-Given the character’s persona $\mathcal{D}_{p}$, ACTS aims to preserve the context between chunks, as well as the information about corresponding subsections within the persona for each chunk, that is, the hierarchical context.
-To this end, ACTS first finds the maximum length of the paragraphs that constitute the persona as follows:
-$l_{\text{max}}=\varphi(p_{1},p_{2},...,p_{l})$ (3)
-, where $\varphi$ denotes a length-calculating function.
-Then, ACTS sets the overlap length of the text splitter to half of $l_{\text{max}}$ (i.e., $l_{\text{o}}=l_{\text{max}}/2$).
-Note that the reason for setting the overlap length sufficiently large is to minimize information loss, as the context between pieces of information contained in each chunk is indispensable in RAG-based role-playing.
-Finally, $\mathcal{D}_{p}$ is segmented using $l_{\text{max}}$ and $l_{\text{o}}$ and hierarchical context $\mathcal{H}_{i}$ is appended to each chunk in order to preserve information such as descriptions of the corresponding characters and situational context at each point in the narrative:
-$\text{ACTS}(\mathcal{D}_{p},\mathcal{H},l_{\text{max}},l_{\text{o}})=\{\hat{c}_{1},\hat{c}_{2},...,\hat{c}_{m}\}$ (4)
-$\hat{c}_{i}=c_{i}+\mathcal{H}_{i}$ (5)
-Guided Selection
+5.2 Evaluation Protocols
+Tasks.
+As we follow the similar experimental protocol in prposed by previous studies (Wang et al., 2024b; Park et al., 2025), we also exploit 60 MBTI questions and 120 BFI (Barrick and Mount, 1991) questions to investigate whether each character can appropriately respond to questions for which they do not have explicit prior knowledge (Figure 4).
+Following the prior work, since it is not possible to construct QA pairs for questions outside the scope of a character’s knowledge, we instead conduct interview-based assessments (Wang et al., 2024b) for each character and compare the results to psychological test outcomes for the character, as determined by thousands of actual participants’ votes [^3] https://www.personality-database.com/(Sang et al., 2022; Wang et al., 2024b).
+We also use 450 QA pairs from the CharacterRAG dataset to verify whether the RPA sufficiently leverages each character’s knowledge.
+Metrics.
+We design three LLM-based metrics, similar to those in prior studies (Wang et al., 2025a, 2024b), to comprehensively evaluate the role-playing capabilities of RAG-based RPAs.: (i) ACC measures whether the character’s response contains the correct answer or not.
+(ii) ACCL is a score assigned by the LLM, ranging from 1 to 10, that evaluates how well the character’s response reflects the correct answer.
+(iii) Hallucination Score (HS) evaluates the degree of hallucination in the model’s response given a query, the relevant chunks, and the ground-truth answer, on a scale from 1 to 10.
+Specifically, HS is assigned close to 1 when the response faithfully reflects only the facts contained in the chunks or answer without distortion or addition, indicating minimal hallucination.
+Table 1: Distribution of Similarity Scores.
+We analyze the distributions of similarity scores to examine whether splitting the text into optimally sized chunks for each character’s persona and incorporating hierarchical context is effective.
+[…]
+$\sum\mu$ ($\uparrow$) $\sum\sigma^{2}$ ($\downarrow$) $\sum\mu$ ($\uparrow$) $\sum\sigma^{2}$ ($\downarrow$) $\sum\mu$ ($\uparrow$) $\sum\sigma^{2}$ ($\downarrow$)
+RCTS 6.4325 0.1026 8.3306 0.1557 12.3136 0.0071
+MHTS 6.4262 0.1038 8.3410 0.1552 12.3063 0.0071
+SC 5.3405 0.1625 8.1691 0.1783 - -
+ATS 6.7007 0.0884 8.4718 0.1281 12.2336 0.0070
+ACTS (Ours) 6.8575 0.0784 8.6226 0.1179 12.3240 0.0063
+Table 2: Human Evaluation. $S$ represents the results of all human evaluators, $\mu$ is $\mathbb{E}(\mathbb{E}(S))$, $\sigma$ is $\mathbb{E}(\sigma(S))$, and Mdn is $\mathbb{E}(Mdn(S)).$
+$\mu$ ($\uparrow$) $\sigma$ ($\downarrow$) Mdn ($\uparrow$) Cronbach’s alpha ($\uparrow$)
+BFI 3.970 0.962 4.217 0.825
+MBTI 3.902 0.915 4.000 0.810
+Table 3: Predicted MBTI Types per Character. Experiments are conducted using GPT-4.1 setting.
+Naive RAG CRAG RAPTOR Adaptive RAG LightRAG AMADEUS (Ours) GT
+Anya Forger ISFP (-2) INTJ (-3) INFJ (-2) INFJ (-2) INFJ (-2) ENFP (0) ENFP
+Chika Fujiwara ENFP (0) ENFP (0) ENFJ (-1) ENFJ (-1) INTP (-2) ENFP (0) ENFP
+Edward Elric INTP (-1) ISTJ (-3) INTP (-1) ENFP (-1) INTP (-1) INFP (-2) ENTP
+Frieren INFP (-1) INFP (-1) INFP (-1) INFP (-1) INTP (0) INTP (0) INTP
+Hitori Gotoh ISFP (-1) ISTJ (-2) INFP (0) INFP (0) ENFP (-1) INFP (0) INFP
+Light Yagami INTJ (-1) INTJ (-1) INTJ (-1) INTJ (-1) INTJ (-1) INTJ (-1) ENTJ
+Māo Māo ISTJ (-2) INTJ (-1) ISTJ (-2) ENTJ (-2) INTJ (-1) ISTP (-1) INTP
+Megumin ISFP (-1) INFP (0) ENFP (-1) ENFP (-1) INFP (0) ISFP (-1) INFP
+Mikoto Misaka ENFP (-2) ISFP (-4) ENFJ (0) ENFJ (0) ENTJ (0) INFJ (-2) ENTJ
+Nina Iseri INFP (-1) ISFP (0) ENFJ (-1) INFJ (-2) ENFJ (-3) INFP (-1) ISFP
+Saitama ISFP (-1) ISTP (0) ISFJ (-2) INFJ (-3) INTP (-1) ISTP (0) ISTP
+Son Goku ESFP (0) ENFJ (-2) ENFJ (-2) ENFP (-1) INTP (-3) ESFP (0) ESFP
+Tanjiro Kamado ENFP (-1) ENFJ (0) ENFJ (0) ENFJ (0) INTP (-3) ENFJ (0) ENFJ
+Tobio Kageyama ENFJ (-3) ENTJ(-2) ISTJ (0) ISTJ (0) INFJ (-1) ISTJ (0) ISTJ
+Yui Hirasawa ISTJ (-4) ENFP (0) ENFP (0) ENFP (0) INTP (-2) ESFP (-1) ENFP
+$\sum|d|$ ($\downarrow$) 21 19 14 15 21 9 -
+Accuracy ($\uparrow$) 65.00% 68.33% 76.67% 75.00% 65.00% 85.00% -
+Avg F1-Score ($\uparrow$) 0.6146 0.6448 0.6925 0.6832 0.5344 0.8244 -
+Table 4: Predicted Big 5 SLOAN Types per Character. Experiments are conducted using GPT-4.1 setting.
+Naive RAG CRAG LightRAG AMADEUS (Ours) GT
+Anya Forger SLOAI (-1) SLOAI (-2) RCUEN (-3) SLUEI (-2) SCUAI
+Chika Fujiwara SCOAI (-1) SCUAI (0) RLUEN (-4) SCOAI (-1) SCUAI
+Edward Elric SCOAI (-3) SLOEI (-1) RCUAN (-4) SLOEI (-1) SLUEI
+Frieren RCOAI (-2) RCUAI (-1) SLUEN (-3) RCUAI (-1) RCUEI
+Hitori Gotoh RLUAI (0) RLUAI (0) RCUEN (-3) RLUAI (0) RLUAI
+Light Yagami SCOEI (-1) SCOEI (-1) RCUAN (-3) SCOEI (-1) RCOEI
+Māo Māo RCOEI (0) RCOAN (-2) SLUAN (-5) RCOEI (0) RCOEI
+Megumin SCOAI (-3) SCUAI (-2) RLUEN (-2) SLOEI (-1) SLUEI
+Mikoto Misaka SLOAI (-3) SLOAI (-3) RCUEN (-2) SLOAI (-3) RCOEI
+Nina Iseri SLOAI (-3) RLUAI (-1) RCUEN (-2) SLUEI (-1) RLUEI
+Saitama RCUAN (0) RCOAN (-1) SCOAI (-3) RCUAN (0) RCUAN
+Son Goku SCOAI (-2) SCUAI (-1) RLUEI (-4) SCOAI (-2) SCUAN
+Tanjiro Kamado SLOAI (-1) SCOAI (0) RCUAN (-3) SCOAI (0) SCOAI
+Tobio Kageyama RCOEN (-1) SLOAN (-2) RCUAI (-4) RCOEN (-1) RLOEN
+Yui Hirasawa SCUAI (0) SLUAI (-1) RCOEN (-4) SCUAI (0) SCUAI
+$\sum|d|$ ($\downarrow$) 21 18 49 14 -
+Accuracy ($\uparrow$) 72.00% 76.00% 34.67% 81.33% -
+Avg F1-Score ($\uparrow$) 0.6785 0.7313 0.2774 0.7986 -
+Table 5: Ablation Study. We analyze the performance gain obtained from incrementally adding each substage.
+Method $\sum|d|$ ($\downarrow$) Accuracy ($\uparrow$) Avg F1-Score ($\uparrow$)
+Naive RAG 21 65.00% 0.6146
+Naive RAG + ACTS 19 68.33% 0.6524
+Naive RAG + ACTS + GS 15 75.00% 0.7426
+Naive RAG + ACTS + GS + AE (Ours) 9 85.00% 0.8394
+Table 6: Role-playing capabilities on CharacterRAG. Higher values of ACC (%) and ACCL (1-10) correspond to better performance, whereas lower values of HS (1-10) are preferable.
+RAG Method GPT-4.1 Gemma3-27B Qwen3-32B
+ACC ($\uparrow$) ACCL ($\uparrow$) HS ($\downarrow$) ACC ($\uparrow$) ACCL ($\uparrow$) HS ($\downarrow$) ACC ($\uparrow$) ACCL ($\uparrow$) HS ($\downarrow$)
+w/o RAG 49.56% 6.79 - 27.56% 5.33 - 18.89% 4.35 -
+Naive RAG 91.33% 9.23 3.13 86.44% 8.85 3.27 78.44% 8.49 5.05
+LightRAG 48.00% 6.06 - 69.56% 8.17 - 68.67% 8.20 -
+CRAG 70.00% 8.26 3.21 57.78% 7.57 4.09 28.67% 5.24 8.68
+AMADEUS (Ours) 92.67% 9.26 2.89 88.00% 8.92 3.26 78.89% 8.63 4.66
+5.3 Experimental Results
+Adaptive Persona Segmentation and Hierarchical Contextualization Are Highly Effective.
+In Table 1, we provide each character with 30 questions, resulting in a total of 450 questions, related to their respective knowledge from the CharacterRAG dataset, and measure the similarity between each question and the chunks retrieved by the RAG model under the three different embedding settings.
+Results demonstrate that compared to RecursiveCharacterTextSplitter (RCTS) (LangChain, 2024b), MarkdownHeaderTextSplitter (MHTS) (LangChain, 2024a), and SemanticChunker (SC) (LangChain, 2024c), adaptive persona segmentation, which we call Adaptive Text Splitter (ATS), segments text with an optimal persona length and overlap for each character and without hierarchical context, achieves a higher average score and lower variance.
+This indicates that each chunk generated using adaptive persona segmentation contains richer semantic information for the same query.
+Building on this, ACTS, which considers hierarchical context in addition to ATS, consistently achieves better performance across all three embedding settings.
+These results show that optimal chunk length, appropriate overlap, and consideration of hierarchical context all play essential roles in effective text chunking.
+Extracting a Character’s Attributes from Selected Chunks Is Reliable.
+We investigate the reasonableness of inferring character attributes with AE from chunks extracted via GS by conducting human evaluation using a 5-point Likert scale.
+To this end, we invite 14 human evaluators and each evaluator is asked to score 60 randomly selected samples.
+Each sample consists of pairs of chunks selected from GS and attributes extracted through AE, for 30 BFI questions and 30 MBTI questions.
+Table 2 shows that the means $\mu$ is close to 4, with small standard deviations $\sigma$.
+It demonstrates that the outputs of GS and AE are reliable and trustworthy, even from a human evaluative perspective.
+We measure Cronbach’s alpha (Cronbach, 1951) to evaluate internal consistency among the human evaluators.
+Note that the Cronbach’s alpha values are 0.825 and 0.810, both exceeding the commonly accepted threshold of 0.7 for acceptable reliability.
+Graph-Based and Web Search-Based RAG Are Unsuitable for RAG-Based RPAs.
+One of the major challenges in retrieval-based role-playing is that, when a RPA receives questions involving knowledge outside a character’s persona, it tends to either overuse irrelevant chunks (Figure 6) or generate uninformative responses (Guo et al., 2024; Shukla et al., 2025; Wang et al., 2025b).
+To investigate whether RAG-based RPAs can handle this problem, we conduct extensive experiments in which we ask 15 characters 60 MBTI questions and 120 BFI questions each, and and evaluate their ability to accurately infer the characters’ personality types.
+Table 3 and Table 4 shows predicted MBTI types and Big 5 SLOAN types per character.
+The number in parentheses indicates the number of times the ground-truth (GT) type of each character was not correctly identified.
+$\sum|d|$ is a measure obtained by summing these values; lower values are preferable.
+Our framework maintains persona consistency even when answering questions that are not explicitly specified in each character’s persona in both MBTI and BFI settings.
+Note that the performance gap of CRAG is significant between the two settings.
+We assume that questions requiring analogical reasoning are difficult to solve even with web search and that the search results may contain non-negligible noise.
+On the other hand, LightRAG exhibits the lowest performance, which shows that graph-based RAG methods are not well suited for RPA applications due to the high cost of graph construction, the difficulty in adding or removing new knowledge, and challenges in maintaining persona consistency.
+While we did not perform a direct comparison, we observed that GraphRAG (Shukla et al., 2025) suffers from similar problems.
+Impact of Each Module in Our Framework.
+Table 5 demonstrates the results of ablation study, evaluated on 60 MBTI questions.
+We find that ACTS, GS, and AE are all highly effective at handling questions that go beyond a character’s knowledge.
+Note that AE achieves the largest performance improvement, highlighting that extracting and incorporating character attributes as input is far more effective than simply providing chunks related to the user query.
+CharacterRAG Dataset Serves as a Valuable Resource for the Construction and Evaluation of RAG-Based RPAs.
+To investigate the factors influencing the performance of role-playing, we conduct a comprehensive interview-based assessments on the generalization capabilities of models with various LLMs and RAG techniques.
+Table 6 and Table 8 present how the ability to accurately answer questions related to the character’s knowledge, which is a core aspect of role-playing, varies across the applied methodologies.
+We first examine to what extent each LLM possesses background knowledge about the 15 characters in a setting without RAG, and results show that none of the three LLMs are capable of effective role-playing without access to external knowledge.
+Moreover, we observe that LightRAG, a graph-based RAG, is ill-suited for the storage and retrieval of character knowledge, as it often suffers from issues such as entity ambiguity and uninformative responses.
+In a similar vein, CRAG exhibits challenges in maintaining role-playing fidelity, which can be attributed to the tendency of web search-based RAG methods to utilize retrieved content that may undermine the consistency of a character’s persona.
+Indeed, despite leveraging web information, CRAG is able to correctly answer only 6 out of the 30 CharacterRAG questions pertaining to Nina Iseri.
+In addition, to analyze how a thinking mode of LLMs influences their role-playing capabilities, we employ Qwen 3-32B.
+Results demonstrate that the thinking mode fails to yield any substantial positive effect on enhancing role-playing performance.
+Our framework achieves the best performance across all three LLMs.
+We also find that the Hallucination Score (HS) is the lowest in CharacterRAG setting.
+6 Conclusion
+In this work, we addresse critical limitations in building retrieval-augmented, RPAs with LLMs.
+By introducing a novel framework consisting of an Adaptive Context-aware Text Splitter (ACTS), Guided Selection (GS), and Attribute Extractor (AE), our approach enables robust and consistent simulation of character personas, even when confronted with queries that extend beyond explicit persona knowledge.
+Through the development of the CharacterRAG dataset, we provide a valuable resource for reproducible evaluation and benchmarking of RAG-based RPAs.
+Our experimental results demonstrate that the proposed method not only enhances the character’s knowledge representation, but also faithfully models nuanced traits such as personality.
+We are enthusiastic about the future prospects of RAG-driven role-playing agents, along with the creation of stronger character personas and improved RAG architectures.
+Ethical Considerations
+This study follows ethical guidelines to ensure the integrity and fairness of all experiments.
+Data and Privacy.
+The Namuwiki portion of our data was gathered within the bounds of the platform’s licensing/usage rules.
+We verified that the collection contained no personally identifiable information, and we restricted its use to academic research.
+For personas and prompting, we relied on either publicly released content, anonymized artifacts, or content created specifically for this study with ethical considerations in mind.
+Experimental Design.
+Our experiments were designed in accordance with rigorous ethical guidelines, with particular emphasis on safeguarding participant data and maintaining privacy.
+To mitigate potential bias, we recruited evaluators with diverse backgrounds and carried out the human evaluation process using procedures intended to be equitable and transparent.
+All participants received clear information regarding the study’s objectives and procedures, and they were allowed to discontinue their participation at any point without negative consequences.
+Through compliance with these practices, we seek to advance AI research in a way that is both scientifically meaningful and ethically grounded, prioritizing privacy, intellectual property rights, and the welfare of everyone involved.
+Appendix A Chunk Duplication Frequencies
+Table 7: Work Genres in Which Each Character Appears. We list the work genres associated with the 15 characters included in the CharacterRAG dataset.
+Character Genre
+Anya Forger Action, Comedy, Drama, Slice of Life, Spy
+Chika Fujiwara Gag, Romantic Comedy
+Edward Elric Adventure, Fantasy
+Frieren Fantasy, Adventure
+Hitori Gotoh Band, Comedy, Drama, Music, Slice of Life
+Light Yagami Crime, Dark Fantasy, Noir, Thriller
+Megumin Comedy, Fantasy, Isekai
+Mikoto Misaka Cyberpunk, SF, Slice of Life
+Nina Iseri Band, Drama
+Saitama Action, Comedy, Hero
+Son Goku Battle, Comedy, Fantasy
+Tanjiro Kamado Action, Adventure, Historical Fantasy
+Tobio Kageyama Sports, Volleyball
+Yui Hirasawa Band, School, Slice of Life
 Algorithm 1 Guided Selection (GS)
-Input: User query $u$; Set of knowledge chunks $\mathcal{C}$; Maximum number of search iterations $N$; Slot size $K$
+Input: User query $u$; Set of knowledge chunks $\mathcal{C}$; Maximum number of search iterations $N$; Slot size $M$
 Output: Selected chunk set $S$
 1: Initialize slot $S=\emptyset$
-2: Sort chunks $\mathcal{C}$ in descending order according to similarity to $u$, obtaining $\mathcal{C}_{\text{sorted}}$
+2: Sort chunks $\mathcal{C}$ in descending order according to semantic similarity to $u$, obtaining $\mathcal{C}_{\text{sorted}}$
 3: Set iteration counter $t\leftarrow 0$
 4:for each chunk $c$ in $\mathcal{C}_{\text{sorted}}$ do
-5:  if $t\geq N$ or $|S|\geq K$ then
+5:  if $t\geq N$ or $|S|\geq M$ then
 6:   break
@@
 14:if $|S|=0$ then
-[…]
+15:  $S\leftarrow$ Top-$K+1$ chunks from $\mathcal{C}_{\text{sorted}}$ (highest semantic similarity to $u$)
 16:end if
 17:return $S$
-While RAG has demonstrated significant potential in improving factual correctness of LLMs, existing RAG methods tend to generate uninformative responses (e.g., I’m sorry, but I don’t have specific information.) for questions outside their knowledge base (Guo et al.
-2024; Shukla et al.
-2025; Wang, Leung, and Shen 2025), or excessively and repetitively use irrelevant chunks that are not pertinent to the given query.
-In this paper, GS focuses on selecting appropriate chunks to generate natural and persona-consistent responses.
-GS is composed of three stages.
-First, we iterate over the chunks, which are sorted in descending order of similarity to the user query $u$, and employ an LLM to determine whether it is possible to infer the corresponding character’s attributes from each chunk for the $u$.
-Second, the chunks selected in the previous step are added to the slot, and the iteration terminates when the slot is full.
-Finally, if the slot remains empty after the maximum number of search iterations, the $K$ chunks with the highest similarity to the query is returned.
-GS is effective in identifying chunks containing information that can be inferred from a character’s actions, such as beliefs or personality traits, which are not explicitly stated in the knowledge base and therefore are difficult to retrieve through direct search.
-For example, if there is no explicit knowledge corresponding to the query ”My living and work spaces are clean and organized” but the information ”Megumin is diligent.” is available, RPA can use this information to infer the characteristics of Megumin and generate an appropriate response.
-GS is summarized in Algorithm 1.
-Attribute Extractor
-Inspired by the observations that incorporating character’s attributes can lead to more realistic responses (Park, Park, and Lim 2025; Chen et al.
-2025), AE considers two attributes: Belief and Value, and Psychological Traits.
-Beliefs and values are fundamental principles and ideological orientations that inform and influence a character’s viewpoints and choices.
-On the other hand, psychological traits refer to characteristics related to personality, emotional states, personal interests, and cognitive tendencies.
-AE extracts attributes of a character from the chunks generated as a result of GS, and dynamically exploits them as context.
-Experiments
-Setup
-Dataset.
-We construct the CharacterRAG dataset, which consists of 15 fictional characters, to leverage a RAG-based role-playing framework.
-CharacterRAG is a high-quality, role-playing dataset in which all external information about works featuring characters that could affect persona consistency has been manually removed, and each persona document has been directly reconstructed from the perspective of each character by human annotators [^1] CharacterRAG dataset is sourced from Namuwiki and is based on Korean data: https://namu.wiki/.
-For instance, any information speculated from the perspective of editors rather than the characters themselves, as well as information such as character popularity polls that may disrupt role-playing, is excluded.
-CharacterRAG consists of 15 distinct fictional characters, 976K written char- acters, and 450 question–answer pairs.
-Six attributes [^2] Unlike the other attributes, Belief and Value and Psychological Traits directly impact the behavior of a character; therefore, AE utilizes information from these two attributes.
-define each character’s persona and the corresponding QA pairs (Chen et al.
-2025):
-  • Activity: A documented history comprising prior activities, behaviors, and interactions, encompassing elements such as backstory and schedules.
-  • Belief and Value: The foundational tenets, dispositions, and ideological orientations that inform and guide a character’s viewpoints and decision-making processes (e.g., beliefs and attitudes).
-  • Demographic Information: Information that can identify a character, including but not limited to their name, age, educational background, professional history, and geographic location.
-  • Psychological Traits: Attributes associated with personality traits, emotional states, preferences, and patterns of cognitive behavior.
-  • Skill and Expertise: The extent of understanding, skillfulness, and competence regarding particular fields or technologies.
-  • Social Relationships: The characteristics and processes of social interactions, encompassing individuals’ roles, relational ties, and patterns of communication.
-[…]
-$\sum\mu$ $\sum\sigma^{2}$ $\sum\mu$ $\sum\sigma^{2}$ $\sum\mu$ $\sum\sigma^{2}$
-Vanilla 6.4325 0.1026 8.3306 0.1557 12.3136 0.0071
-ATS 6.7007 0.0884 8.4718 0.1281 12.2336 0.0070
-ACTS (Ours) 6.8575 0.0784 8.6226 0.1179 12.3240 0.0063
-Table 1: Distribution of Similarity Scores.
-Vanilla refers to the method of splitting chunks with a fixed length and overlap, ATS refers to dividing a persona for each character with optimal segment length and overlap, without hierarchical context.
-$\mu$ $\sigma$ Mdn Cronbach’s alpha
-BFI 3.890 0.965 4.100 0.845
-MBTI 3.870 0.911 3.933 0.830
-Table 2: Human Evaluation.
-We conduct human evaluation using a 5-point Likert scale to verify whether inferring character attributes with the Attribute Extractor (AE) from chunks extracted by Guided Selection (GS) is reasonable.
-Note that $S$ represents the results of all human evaluators, $\mu$ is $\mathbb{E}(\mathbb{E}(S))$, $\sigma$ is $\mathbb{E}(\sigma(S))$, and Mdn is $\mathbb{E}(Mdn(S)).$
-Chunk Usage Rate GS Failure
-CharacterRAG 67.56% 25.11%
-MBTI 43.84% 15.33%
-Table 3: Average Chunk Usage Rate and GS Failure. We present 30 CharacterRAG questions and 60 MBTI questions to 15 characters, and measure the average chunk usage rate and GS failure.
-Each section of the character’s persona contains subsections, preserving the hierarchical context of each chunk (e.g., ”Frieren’s activities in the story, the journey to find Aureole, the First-Class Mage Exam”).
-Furthermore, Each QA pair consists of a question and corresponding answer derived from the character’s knowledge, pertaining to one of the six attributes manually constructed for each character.
-Detailed statistics of CharacterRAG are shown in Figure 3.
-[…]
-Settings.
-We implement Guided Selection (GS) and Attribute Extractor (AE) using GPT-4.1 (”gpt-4.1-2025-04-14”).
-We apply the ACTS to the conventional Naive RAG.
-The maximum number of search iterations $N$ is 30, and the slot size $K$ is set to 2.
-Note that the speed at which GS identifies and retrieves relevant chunks for answering is much faster and more cost-efficient than that of previous graph-based or web search-based RAG methods (Shukla et al.
-2025; Guo et al.
-2024; Yan et al.
-2024).
-Evaluation Protocols
-Tasks.
-We use 450 QA pairs from the CharacterRAG dataset to verify whether the RPA sufficiently leverages each character’s knowledge.
-We also use 60 MBTI [^3] https://www.16personalities.com/ questions and 120 BFI (Barrick and Mount 1991) questions to investigate whether each character can appropriately respond to questions for which they do not have explicit prior knowledge.
-Since it is not possible to construct QA pairs for questions outside the scope of a character’s knowledge, we instead conduct an interview-based assessment (Wang et al.
-2024b) for each character and compare the results to psychological test outcomes for the character, as determined by thousands of actual participants’ votes [^4] https://www.personality-database.com/.
-Metrics.
-We design three metrics to extensively evaluate the role-playing capabilities of RAG-based RPAs: (i) ACC measures whether the character’s response contains the correct answer or not.
-(ii) ACCL is a score assigned by the LLM, ranging from 1 to 10, that evaluates how well the character’s response reflects the correct answer.
-(iii) Hallucination Score (HS) evaluates the degree of hallucination in the model’s response given a query, the relevant chunks, and the ground-truth answer, on a scale from 1 to 10.
-Specifically, HS is assigned close to 1 when the response faithfully reflects only the facts contained in the chunks or answer without distortion or addition, indicating minimal hallucination.
-CRAG LightRAG AMADEUS (Ours) GT
-Anya Forger INTJ (-3) INFJ (-2) ENFP (0) ENFP
-Chika Fujiwara ENFP (0) INTP (-2) ENFP (0) ENFP
-Edward Elric ISTJ (-3) INTP (-1) INFP (-2) ENTP
-Frieren INFP (-1) INTP (0) INTP (0) INTP
-Hitori Gotoh ISTJ (-2) ENFP (-1) INFP (0) INFP
-Light Yagami INTJ (-1) INTJ (-1) INTJ (-1) ENTJ
-Māo Māo INTJ (-1) INTJ (-1) ISTP (-1) INTP
-Megumin INFP (0) INFP (0) ISFP (-1) INFP
-Mikoto Misaka ISFP (-4) ENTJ (0) INFJ (-2) ENTJ
-Nina Iseri ISFP (0) ENFJ (-3) INFP (-1) ISFP
-Saitama ISTP (0) INTP (-1) ISTP (0) ISTP
-Son Goku ENFJ (-2) INTP (-3) ESFP (0) ESFP
-Tanjiro Kamado ENFJ (0) INTP (-3) ENFJ (0) ENFJ
-Tobio Kageyama ENTJ(-2) INFJ (-1) ISTJ (0) ISTJ
-Yui Hirasawa ENFP (0) INTP (-2) ESFP (-1) ENFP
-Accuracy 68.33% 65.00% 85.00% -
-Avg F1-Score 0.6448 0.5344 0.8244 -
-Table 4: Predicted MBTI Types per Character.
-The number in parentheses indicates the number of times the ground-truth (GT) type of each character was not correctly identified (using the GPT-4.1 setting).
-Experimental Results
-[…]
-CRAG LightRAG AMADEUS (Ours) GT
-Anya Forger SLOAI (-2) RCUEN (-3) SLUEI (-2) SCUAI
-Chika Fujiwara SCUAI (0) RLUEN (-4) SCOAI (-1) SCUAI
-Edward Elric SLOEI (-1) RCUAN (-4) SLOEI (-1) SLUEI
-Frieren RCUAI (-1) SLUEN (-3) RCUAI (-1) RCUEI
-Hitori Gotoh RLUAI (0) RCUEN (-3) RLUAI (0) RLUAI
-Light Yagami SCOEI (-1) RCUAN (-3) SCOEI (-1) RCOEI
-Māo Māo RCOAN (-2) SLUAN (-5) RCOEI (0) RCOEI
-Megumin SCUAI (-2) RLUEN (-2) SLOEI (-1) SLUEI
-Mikoto Misaka SLOAI (-3) RCUEN (-2) SLOAI (-3) RCOEI
-Nina Iseri RLUAI (-1) RCUEN (-2) SLUEI (-1) RLUEI
-Saitama RCOAN (-1) SCOAI (-3) RCUAN (0) RCUAN
-Son Goku SCUAI (-1) RLUEI (-4) SCOAI (-2) SCUAN
-Tanjiro Kamado SCOAI (0) RCUAN (-3) SCOAI (0) SCOAI
-Tobio Kageyama SLOAN (-2) RCUAI (-4) RCOEN (-1) RLOEN
-Yui Hirasawa SLUAI (-1) RCOEN (-4) SCUAI (0) SCUAI
-Accuracy 76.00% 34.67% 81.33% -
-Avg F1-Score 0.7313 0.2774 0.7986 -
-Table 5: Predicted Big 5 SLOAN Types per Character.
-The number in parentheses indicates the number of times the ground-truth (GT) type of each character was not correctly identified (using the GPT-4.1 setting).
-Extracting a character’s attributes from selected text chunks is reliable.
-We investigate the reasonableness of inferring character attributes with the Attribute Extractor (AE) from chunks extracted via Guided Selection (GS) by conducting human evaluation using a 5-point Likert scale.
-To this end, we invite 11 human evaluators and each evaluator is asked to score 60 randomly selected samples.
-Each sample consists of pairs of chunks selected from GS and attributes extracted through AE, for 30 BFI questions and 30 MBTI questions that are not included in the character’s knowledge.
-In Table 2, we find that the means $\mu$ is close to 4, with small standard deviations $\sigma$.
-It demonstrates that the outputs of GS and AE are reliable and trustworthy, even from a human evaluative perspective.
-We also measure Cronbach’s alpha (Cronbach 1951) to evaluate internal consistency among the human evaluators.
-We find that the Cronbach’s alpha values are 0.845 and 0.830, both exceeding the commonly accepted threshold of 0.7 for acceptable reliability.
-Since values above 0.8 are generally interpreted as indicating a high level of internal consistency, experimental results in Table 2 can be considered highly trustworthy.
-Selecting appropriate chunks increases the utilization of a character’s persona.
-Table 3 shows the usage rate of persona chunks and the failure rate of GS when each each of the 15 characters is presented with questions that are both within and outside their knowledge.
-Note that, despite the limited number of queries in CharacterRAG, the Chunk Usage Rate still reaches 67.56%.
-The findings demonstrate that GS serves as an effective tool for enabling the use of character personas.
-At the same time, it can be observed that the 450 questions in CharacterRAG are designed to capture the core information of each character.
-Meanwhile, as shown in Table 3 and Figure 1, GS substantially increases the chunk utilization rate for the 60 MBTI questions, from 34.93% to 43.84%.
-The proportion of cases where the GS fails to find an appropriate chunk, i.e., GS Failure, is 25.11% and 15.33%, respectively, which demonstrates that although GS effectively utilizes the character’s persona, there is still room for further improvement.
-Nevertheless, despite this shortcoming, our framework offers markedly enhanced speed, cost-effectiveness, and stable performance compared to earlier RAG approaches utilizing graph structures or web search mechanisms.
-(Shukla et al.
-2025; Guo et al.
-2024; Yan et al.
-2024).
-Method Task GPT-4.1 Gemma3-27B Qwen3-32B
-[…]
-w/o RAG CharacterRAG 49.56% 6.79 - 27.56% 5.33 - 18.89% 4.35 -
-Naive RAG CharacterRAG 91.33% 9.23 3.13 86.44% 8.85 3.27 78.44% 8.49 5.05
-LightRAG CharacterRAG 48.00% 6.06 - 69.56% 8.17 - 68.67% 8.20 -
-CRAG CharacterRAG 70.00% 8.26 3.21 57.78% 7.57 4.09 28.67% 5.24 8.68
-AMADEUS (Ours) CharacterRAG 92.67% 9.26 2.89 88.00% 8.92 3.26 78.89% 8.63 4.66
-Naive RAG MBTI - - 2.69 - - 2.53 - - 2.33
-CRAG MBTI - - 2.38 - - 2.91 - - 1.80
-AMADEUS (Ours) MBTI - - 2.05 - - 2.02 - - 2.04
-Naive RAG BFI - - 2.74 - - 2.52 - - 2.42
-CRAG BFI - - 2.26 - - 2.75 - - 1.96
-AMADEUS (Ours) BFI - - 1.94 - - 1.99 - - 2.03
-Table 6: Evaluation of role-playing capabilities on CharacterRAG, MBTI, and BFI tasks.
-Higher values of ACC (%) and ACCL (1-10) correspond to better performance, whereas lower values of HS (1-10) are preferable.
-Graph-based RAG and web search-based RAG are unsuitable for role-playing.
-One of the major challenges in retrieval-based role-playing is that, when a RPA receives questions involving knowledge outside a character’s persona, it tends to either overuse irrelevant chunks (Figure 1) or generate uninformative responses (Guo et al.
-2024; Shukla et al.
-2025; Wang, Leung, and Shen 2025).
-To investigate whether RAG-based RPAs can handle this problem, we conduct extensive experiments in which we ask 15 characters 60 MBTI questions and 120 BFI questions each, and and evaluate their ability to accurately infer the characters’ personality types.
-Table 4 and Table 5 show predicted MBTI types and Big 5 SLOAN types per character.
-Our framework maintains persona consistency even when answering questions that are not explicitly specified in each character’s persona in both MBTI and BFI settings.
-Note that the performance gap of CRAG is significant between the two settings.
-We assume that questions requiring analogical reasoning are difficult to solve even with web search and that the search results may contain non-negligible noise.
-On the other hand, LightRAG exhibits the lowest performance, which shows that graph-based RAG methods are not well suited for RPA applications due to the high cost of graph construction, the difficulty in adding or removing new knowledge, and challenges in maintaining persona consistency.
-While we did not perform a direct comparative experiment, we observed that GraphRAG (Shukla et al.
-2025) suffers from similar problems.
-Specifically, applying graph-based RAG to RPA is challenging in cases where the persona frequently changes, such as with daily streamers or VTubers.
-CharacterRAG dataset serves as a valuable resource for the construction and evaluation of RAG-based RPAs.
-To investigate the factors influencing the performance of role-playing, we conduct a comprehensive interview-based assessments on the generalization capabilities of models with various LLMs and RAG techniques.
-Table 6 presents how the ability to accurately answer questions related to the character’s knowledge, which is a core aspect of role-playing, varies across the applied methodologies.
-We first examine to what extent each LLM possesses background knowledge about the 15 characters in a setting without RAG, and results show that none of the three LLMs are capable of effective role-playing without access to external knowledge.
-Moreover, we observe that LightRAG, a graph-based RAG, is ill-suited for the storage and retrieval of character knowledge, as it often suffers from issues such as entity ambiguity and uninformative responses.
-In a similar vein, CRAG exhibits challenges in maintaining role-playing fidelity, which can be attributed to the tendency of web search-based RAG methods to utilize retrieved content that may undermine the consistency of a character’s persona.
-Indeed, despite leveraging web information, CRAG is able to correctly answer only 6 out of the 30 CharacterRAG questions pertaining to Nina Iseri.
-In addition, to analyze how a thinking mode of LLMs influences their role-playing capabilities, we employ Qwen 3-32B.
-Results demonstrate that the thinking mode fails to yield any substantial positive effect on enhancing role-playing performance.
-Note that our framework achieves the best performance across all three LLMs.
-We also find that the Hallucination Score (HS) is the lowest in CharacterRAG setting.
-These results highlight the importance of preserving the context of split-character personas and effectively leveraging appropriate character attributes in RAG-based RPAs.
-Furthermore, such elements are especially pronounced in dialogue situations that transcend the scope of the character’s knowledge (Table 4 and Table 5).
-We believe that our findings demonstrate new possibilities for RAG-based RPAs.
-Conclusion
-In this work, we addresse critical limitations in building retrieval-augmented, RPAs with LLMs.
-By introducing a novel framework consisting of an Adaptive Context-aware Text Splitter (ACTS), Guided Selection (GS), and Attribute Extractor (AE), our approach enables robust and consistent simulation of character personas, even when confronted with queries that extend beyond explicit persona knowledge.
-Through the development of the CharacterRAG dataset, we provide a valuable resource for reproducible evaluation and benchmarking of RAG-based RPAs.
-Our experimental results demonstrate that the proposed method not only enhances the character’s knowledge representation, but also faithfully models nuanced traits such as personality.
-We are enthusiastic about the future prospects of RAG-driven role-playing agents, along with the creation of stronger character personas and improved RAG architectures.
+In Figure 6, previous RAG method retrieves (k = 3) relevant chunks for each of the 60 given MBTI questions, and subsequently measures both the chunk utilization rate and the duplication frequency.
+Our method measures the chunk utilization rate and duplication frequency for the chunks selected by ACTS and GS.
+Appendix B Genre Diversity in CharacterRAG
+Table 7 shows the genres of works in which each character appears.
+Note that while there are some similar categories, CharacterRAG encompasses a broader range of genres.
+The detailed personality distribution for each character (e.g., Extroverted vs.
+Introverted) is shown in Figure 2.
+Appendix C Related Work
+Role‑Playing Agents.
+With the advent of LLMs, researchers have pursued finer‑grained persona consistency (Zhang et al., 2018; Ji et al., 2025; Park et al., 2025; Lu et al., 2024; Zhou et al., 2024).
+Complementary benchmarks soon followed, along with various evaluation methods (Boudouri et al., 2025; Ahn et al., 2024; Wang et al., 2024b, 2025a).
+However, there has been little research on RAG-based role-playing agents (RPAs).
+In this paper, we propose AMADEUS, a RAG-based RPA framework that not only elicits information related to a character, but also maintains persona consistency even when responding to queries beyond its explicit knowledge.
+Retrieval‑Augmented Generation (RAG).
+RAG couples non-parametric memory with an LLM to mitigate hallucination and stale knowledge (Lewis et al., 2020; Bhat et al., 2025; Zhong et al., 2025b; Zhu et al., 2024).
+Nevertheless, no existing benchmark explicitly targets role‑playing under RAG, and prior work still assumes personas are short, knowledge‑dense snippets.
+We mitigate this discrepancy with a novel text splitter that tailors chunk lengths and hierarchical context to each character.
+We also introduce CharacterRAG, the first dataset designed for building and evaluating RAG‑based role‑playing agents across 15 fictional character’s personas.
+Table 8: Role-playing capabilities on MBTI and BFI. Lower values of HS (1-10) are preferable.
+RAG Method GPT-4.1 Gemma3-27B Qwen3-32B
+HS ($\downarrow$) HS ($\downarrow)$ HS ($\downarrow$)
+MBTI
+Naive RAG 2.69 2.53 2.33
+CRAG 2.38 2.91 1.80
+AMADEUS (Ours) 2.05 2.02 2.04
+BFI
+Naive RAG 2.74 2.52 2.42
+CRAG 2.26 2.75 1.96
+AMADEUS (Ours) 1.94 1.99 2.03
+Prompt used in Guided Selection In the given [persona_name]’s Context, output True if you can clearly infer the Belief and Value information or Psychological Traits of [persona_name] to answer the given User Query; otherwise, output False.
+Output only True or False.
+# Explanation of Belief and Value:
+- The principles, attitudes, and ideological stances that shape an individual’s perspectives and decisions
+- e.g.,
+[persona_name]’s "beliefs", "attitudes", "values", "religion"
+# Explanation of Psychological Traits:
+- Characteristics related to personality, emotions, interests, and cognitive tendencies.
+- e.g.,
+[persona_name]’s "Personality", "hobby and interest", "emotional"
+Figure 3: Prompt used in Guided Selection (GS). GS returns whether Belief and Value information or Psychological Traits can be inferred from given chunks.
+Prompt ued in the interview-based psychological assessment (MBTI) Based on the given [persona_name]’s answer, please rate the question, "Am I ([persona_name]) [question]?" on a scale from 1 to 7.
+A score closer to 1 means "very much so," and a score closer to 7 means "not at all."
+Output only the number.
+Figure 4: Prompt ued in the interview-based psychological assessment (MBTI). In the BFI test, responses are rated on a scale from 1 to 5.
+Figure 5: AMADEUS framework.
+AMADEUS consists of three substage: (i) ACTS splits a persona document to make it suitable for RAG-based role-playing.
+(ii) To fully leverage the character’s knowledge, GS retrieves chunks from which it can infer the answer to the user query.
+(iii) AE uses the information derived from the GS results to extract character attributes.
+Figure 6: Chunk Duplication Frequencies.
+We compare the distribution of chunk duplication frequencies and chunk usage rates between Naive RAG and our method when questions involving knowledge not present in the persona document were given.
+We observe that when each of the 60 MBTI questions is asked to 15 characters, the average chunk usage rate increases from $34.93\%$ to $43.84\%$, and the distribution becomes more uniform.
+Figure 7: An example of the output produced at each substage of Amadeous.
+To improve interpretability, we provide a detailed breakdown of the results for a character from the CharacterRAG dataset (Yui Hirasawa).
+Appendix D Settings
+We implement GS and AE using GPT-4.1 ("gpt-4.1-2025-04-14").
+We develop the Amadeus framework based on naïve RAG.
+The maximum number of search iterations $N$ is 30, and the slot size $M$ is set to 2.
+We performed benchmarking on an NVIDIA H100 NVL GPU.
+Appendix E Log-Density Ridgeline Analysis
+To empirically validate the suitability of the overlap coefficient, Figure 8 presents the log-density ridgelines of five distributions estimated under the normality assumption: $\log f\!\left(x\mid\sum\mu,\,\sum\sigma^{2}\right)=-0.5\left(\frac{x-\sum\mu}{\sqrt{\sum\sigma^{2}}}\right)^{2}-\log\!\left(\sqrt{\sum\sigma^{2}}\right)-0.5\log(2\pi)$.
+We observe that, when $\alpha=2$, the sum of the similarity scores is maximized while their variance is minimized.
+Figure 8: Ridgeline of Log-Density Comparison. Empirical verification of optimal overlap coefficient $\alpha$: $l_{\text{o}}=l_{\text{max}}/\alpha$ (Normal assumption).

### v4

@@
 6 Conclusion
-In this work, we addresse critical limitations in building retrieval-augmented, RPAs with LLMs.
+In this work, we address critical limitations in building retrieval-augmented, RPAs with LLMs.
 By introducing a novel framework consisting of an Adaptive Context-aware Text Splitter (ACTS), Guided Selection (GS), and Attribute Extractor (AE), our approach enables robust and consistent simulation of character personas, even when confronted with queries that extend beyond explicit persona knowledge.
-Through the development of the CharacterRAG dataset, we provide a valuable resource for reproducible evaluation and benchmarking of RAG-based RPAs.
-Our experimental results demonstrate that the proposed method not only enhances the character’s knowledge representation, but also faithfully models nuanced traits such as personality.
-We are enthusiastic about the future prospects of RAG-driven role-playing agents, along with the creation of stronger character personas and improved RAG architectures.
+Through the development of the CharacterRAG dataset, we provide a valuable resource that enables the construction of RAG-based RPAs, as well as their reproducible evaluation.
+Experimental results demonstrate that the proposed method not only enhances the character’s knowledge representation, but also faithfully models nuanced traits such as personality.
+We are enthusiastic about the future prospects of RAG-driven role-playing agents, along with the creation of stronger character persona dataset and improved RAG architectures.
 Ethical Considerations
-This study follows ethical guidelines to ensure the integrity and fairness of all experiments.
+This work follows ethical guidelines to ensure the integrity and fairness of all experiments.
 Data and Privacy.
@@
 Figure 5: AMADEUS framework.
-AMADEUS consists of three substage: (i) ACTS splits a persona document to make it suitable for RAG-based role-playing.
+AMADEUS consists of three substages: (i) ACTS splits a persona document to make it suitable for RAG-based role-playing.
 (ii) To fully leverage the character’s knowledge, GS retrieves chunks from which it can infer the answer to the user query.